soybean meal is 12% protein; cornmeal is 6% protein. How many pound of each should be mixed together in order to get a 240-lb mixture that is 9% protein?
Answer by Boreal(15235)

soybean =x
cornmeal is 240-x
.12x+ .06 (240-x)= .09(240)
.12x + 14.4-.06x=21.6
.06x=7.2
multiply by 100
6x=720
x=120 pounds of soybean
240-x=120 pounds of cornmeal
This makes sense, since 9% is exactly in the middle between 6% of one and 12% of the other.